OpenStudy (anonymous):

Rationalize the denominator in the expression. sq. root of 12x^7 over sq. root of y^9

OpenStudy (anonymous):

there was no reason for issuing warning letter

OpenStudy (anonymous):

first off \(\sqrt{y^9}=\sqrt{y^8}\sqrt{y}=y^4\sqrt{y}\) and similarly \[\sqrt{12x^7}=\sqrt{4x^6}\sqrt{3x}=2x^3\sqrt{3x}\] so you should start with \[\frac{2x^3\sqrt{3y}}{y^4\sqrt{y}}\]

OpenStudy (anonymous):

oops typo there, should be \[\frac{2x^3\sqrt{3x}}{y^4\sqrt{y}}\]

OpenStudy (anonymous):

you are making this more and more complicated

OpenStudy (anonymous):